# 00footnotetext: Shared first author; Shared last author

### v1

@@ Appendix I Model-based quality filters
        [WET file]              A Guide To Markets - The New York Times        {internallinenumbers*} NYTimes.com no longer supports Internet Explorer 9 or earlier. Please upgrade your browser. LEARN MORE »        Sections        Home        Search        Skip to content Skip to navigation View mobile version        The New York Times        Archives|A Guide To Markets        Search        Subscribe Now        Log In        0        Settings        Close search        Site Search Navigation        Search NYTimes.com        Clear this text input        Go        https://nyti.ms/29nVV3Q        Loading...        See next articles        See previous articles        Site Navigation        Site Mobile Navigation        Advertisement        Archives | 1990        A Guide To Markets        MAY 10, 1990        Continue reading the main story Share This Page        Continue reading the main story        About the Archive        {internallinenumbers*} This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.        {internallinenumbers*} Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.        May 10, 1990, Page 00006 The New York Times Archives        {internallinenumbers*} HERE is a sampling of some of the better antiques and flea markets around the United States.        Two or Three Times a Year        {internallinenumbers*} BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.        {internallinenumbers*} RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.        {internallinenumbers*} FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.        Advertisement        Continue reading the main story        Monthly        {internallinenumbers*} ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.        Continue reading the main story        Advertisement        Continue reading the main story        {internallinenumbers*} KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.        {internallinenumbers*} THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.        {internallinenumbers*} SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.        Weekly        {internallinenumbers*} BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.        {internallinenumbers*} LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.        {internallinenumbers*} ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.        {internallinenumbers*} A version of this list appears in print on May 10, 1990, on Page C00006 of the National edition with the headline: A Guide To Markets. Order Reprints| Today’s Paper|Subscribe        Continue reading the main story        What’s Next        Loading...        Go to Home Page »        Site Index The New York Times        Site Index Navigation        News        World        U.S.        Politics        N.Y.        Business        Tech        Science        Health        Sports        Education        Obituaries        Today’s Paper        Corrections        Opinion        Today’s Opinion        Op-Ed Columnists        Editorials        Op-Ed Contributors        Letters        Sunday Review        Video: Opinion        Arts        Today’s Arts        Art & Design        Books        Dance        Movies        Music        N.Y.C. Events Guide        Television        Theater        Video: Arts        Living        Automobiles        Crossword        Food        Education        Fashion & Style        Health        Jobs        Magazine        N.Y.C. Events Guide        Real Estate        T Magazine        Travel        Weddings & Celebrations        Listings & More        Reader Center        Classifieds        Tools & Services        N.Y.C. Events Guide        Multimedia        Photography        Video        NYT Store        Times Journeys        Subscribe        Manage My Account        NYTCo        Subscribe        Subscribe        Home Delivery        Digital Subscriptions        Crossword        Email Newsletters        Gift Subscriptions        Group Subscriptions        Education Rate        Mobile Applications        Replica Edition        Site Information Navigation        © 2019 The New York Times Company        Home        Search        {internallinenumbers*} Accessibility concerns? Email us at accessibility@nytimes.com. We would love to hear from you.        Contact Us        Work With Us        Advertise        Your Ad Choices        Privacy        Terms of Service        Terms of Sale        Site Information Navigation        Site Map        Help        Site Feedback        Subscriptions
@@
        [Trafilatura]              Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.        {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.        So when do you use an en-dash? Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”        What about an em-dash? From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.        {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
@@
        [Resiliparse]              1              Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?              When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?              {internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32              {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.              0              {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:              {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.              So when do you use an en-dash? Again from Mental Floss:              {internallinenumbers*} 1. To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)              {internallinenumbers*} 2. The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”              What about an em-dash? From here:              {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.              {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.              {internallinenumbers*} Not the answer you’re looking for? Browse other questions tagged or ask your own question.
@@
        [WET file]              {internallinenumbers*} syntax - What’s the difference between - and -- in a phrase? - English Language & Usage Stack Exchange        Stack Exchange Network        {internallinenumbers*} Stack Exchange network consists of 175 Q&A communities including Stack Overflow, the largest, most trusted online community for developers to learn, share their knowledge, and build their careers.        Visit Stack Exchange        Log In Sign Up        current community        English Language & Usage        help chat        English Language & Usage Meta        your communities        Sign up or log in to customize your list.        more stack exchange communities        company blog        Tour Start here for a quick overview of the site        Help Center Detailed answers to any questions you might have        Meta Discuss the workings and policies of this site        About Us Learn more about Stack Overflow the company        Business Learn more about hiring developers or posting ads with us        {internallinenumbers*} By using our site, you acknowledge that you have read and understand our Cookie Policy, Privacy Policy, and our Terms of Service.        {internallinenumbers*} English Language & Usage Stack Exchange is a question and answer site for linguists, etymologists, and serious English language enthusiasts. Join them; it only takes a minute:        Sign up        Here’s how it works:        Anybody can ask a question        Anybody can answer        The best answers are voted up and rise to the top        Home        Questions        Tags        Users        Unanswered        What’s the difference between - and — in a phrase? [duplicate]        Ask Question        1        Possible Duplicate:        When should I use an em-dash, an en-dash, and a hyphen?        When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?        syntax dashes symbols        share|improve this question        edited Apr 13 ’17 at 12:38        Community        1        asked Jul 13 ’11 at 0:10        curiouscurious        123115        {internallinenumbers*} marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32        {internallinenumbers*} This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.        add a comment |        1 Answer 1        active oldest votes        0        {internallinenumbers*} The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:        {internallinenumbers*} An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.        So when do you use an en-dash? Again from Mental Floss:        {internallinenumbers*} To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)        {internallinenumbers*} The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”        What about an em-dash? From here:        {internallinenumbers*} Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.        {internallinenumbers*} If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.        share|improve this answer        edited Aug 4 ’15 at 17:00        zwol        2,51911424        answered Jul 13 ’11 at 0:16        simchonasimchona        30.9k5112139        add a comment |        {internallinenumbers*} Not the answer you’re looking for? Browse other questions tagged syntax dashes symbols or ask your own question.        asked        7 years, 9 months ago        viewed        2,336 times        active        3 years, 8 months ago        Featured on Meta        Announcing the arrival of Valued Associate #679: Cesar Manara        Planned maintenance scheduled April 17/18, 2019 at 00:00UTC (8:00pm US/Eastern)        Linked        279        When should I use an em-dash, an en-dash, and a hyphen?        Related        279        When should I use an em-dash, an en-dash, and a hyphen?        4        When to use -, – and —?        6        What is the difference between `-` and `--`        5        Does “cost-benefit ratio” use a hyphen or an en-dash?        0        {internallinenumbers*} What kind of dash character should I use at the end of a famous saying to mark of the author?        -1        dash non-restrictive element in the middle of a sentence        1        em dash followed by a comma        0        what’s the difference between a hyphen, a dash and a minus sign?        0        Using comma to delimit the name of a group and its constituents?        1        I tend to overuse the hyphen as a pause, and would appreciate some feedback on this        Hot Network Questions        How could we fake a moon landing now?        Using audio cues to encourage good posture        Chinese Seal on silk painting - what does it mean?        How does the math work when buying airline miles?        What causes the direction of lightning flashes?        Maximum summed subsequences with non-adjacent items        Most bit efficient text communication method?        Significance of Cersei’s obsession with elephants?        Is it a good idea to use CNN to classify 1D signal?        What font is "z" in "z-score"?        How can I use the Python library networkx from Mathematica?        Fundamental Solution of the Pell Equation        Dating a Former Employee        Do I really need recursive chmod to restrict access to a folder?        Generate an RGB colour grid        How to Make a Beautiful Stacked 3D Plot        How to find all the available tools in mac terminal?        {internallinenumbers*} Has negative voting ever been officially implemented in elections, or seriously proposed, or even studied?        What is the meaning of the simile “quick as silk”?        Circuit to "zoom in" on mV fluctuations of a DC signal?        {internallinenumbers*} What are the out-of-universe reasons for the references to Toby Maguire-era Spider-Man in Into the Spider-Verse?        What does this Jacques Hadamard quote mean?        Can a new player join a group only when a new campaign starts?        Why wasn’t DOSKEY integrated with COMMAND.COM?        more hot questions        English Language & Usage        Tour        Help        Chat        Contact        Feedback        Mobile        Company        Stack Overflow        Stack Overflow Business        Developer Jobs        About        Press        Legal        Privacy Policy        Stack Exchange        Network        Technology        Life / Arts        Culture / Recreation        Science        Other        Stack Overflow        Server Fault        Super User        Web Applications        Ask Ubuntu        Webmasters        Game Development        TeX - LaTeX        Software Engineering        Unix & Linux        Ask Different (Apple)        WordPress Development        Geographic Information Systems        Electrical Engineering        Android Enthusiasts        Information Security        Database Administrators        Drupal Answers        SharePoint        User Experience        Mathematica        Salesforce        ExpressionEngine® Answers        Stack Overflow em Português        Blender        Network Engineering        Cryptography        Code Review        Magento        Software Recommendations        Signal Processing        Emacs        Raspberry Pi        Stack Overflow Ha pycckom        Programming Puzzles & Code Golf        Stack Overflow en español        Ethereum        Data Science        Arduino        Bitcoin        more (31)        Photography        Science Fiction & Fantasy        Graphic Design        Movies & TV        Music: Practice & Theory        Worldbuilding        Seasoned Advice (cooking)        Home Improvement        Personal Finance & Money        Academia        Law        more (15)        English Language & Usage        Skeptics        Mi Yodeya (Judaism)        Travel        Christianity        English Language Learners        Japanese Language        Arqade (gaming)        Bicycles        Role-playing Games        Anime & Manga        Puzzling        Motor Vehicle Maintenance & Repair        more (33)        MathOverflow        Mathematics        Cross Validated (stats)        Theoretical Computer Science        Physics        Chemistry        Biology        Computer Science        Philosophy        more (10)        Meta Stack Exchange        Stack Apps        API        Data        Blog        Facebook        Twitter        LinkedIn        {internallinenumbers*} site design / logo © 2019 Stack Exchange Inc; user contributions licensed under cc by-sa 3.0 with attribution required. rev 2019.4.18.33353        English Language & Usage Stack Exchange works best with JavaScript enabled
@@
MinHash is a locality-sensitive hashing technique used to group sets into collections based on their Jaccard similarity [26]. In the context of deduplicating text datasets, MinHash was first employed in Lee et al. [86] and then used in numerous other projects [118, 41]. We point readers to the main text of Lee et al. [86] and Appendix G.3.1 of Penedo et al. [118] for more details. The primary hyperparameters of note are the ngram-size, and the number of permutations used. Following Lee et al. [86], Penedo et al. [118], we use an ngram-size of 5 tokens and target a Jaccard similarity of 0.8. Departing from prior work, however, we modify the number of MinHash permutations used. Both Lee et al. [86] and Penedo et al. [118] use a total of 9,000 permutations, split into 450 buckets of 20 hashes each. We found this to be overly expensive and notice that similar Jaccard similarity plots can be attained with a much smaller number of permutations. For all of our ablations, we instead use a total of 1,395 permutations, split into 93 buckets of size 15. These hyperparameters were chose programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget. See Figure˜6 for more details.
@@
    In choosing a false positive rate for the ngram-based Bloom filter, it’s important to recognize that removal of a paragraph or document is dictated by having greater than a threshold fraction of the ngrams contained in the set. As an example, suppose we are given a paragraph of $N$ ngrams, where $S$ of them are already contained in the Bloom filter and we set threshold to $T$. Because Bloom filters do not allow false negatives, every one of the $S$ ngrams are marked (correctly) as contained, and NSN-S of them could potentially be marked as a false positive. Indeed, of the NSN-S of these ngrams, at least TNSTN-S of them would need to be marked as a false positive, each of which occurs independently with probability $\epsilon$. This is equivalent to NSN-S Bernoulli random variables with parameter $\epsilon$, and can be bounded by a crude Hoeffding bound. In this particular case, the probability that a document or paragraph is falsely marked as a duplicate is bounded by:

### v2

@@ Appendix I Model-based quality filters
     [WET file]    A Guide To Markets - The New York Times  NYTimes.com no longer supports Internet Explorer 9 or earlier. Please upgrade your browser. LEARN MORE »  Sections  Home  Search  Skip to content Skip to navigation View mobile version  The New York Times  Archives|A Guide To Markets  Search  Subscribe Now  Log In  0  Settings  Close search  Site Search Navigation  Search NYTimes.com  Clear this text input  Go  https://nyti.ms/29nVV3Q  Loading...  See next articles  See previous articles  Site Navigation  Site Mobile Navigation  Advertisement  Archives | 1990  A Guide To Markets  MAY 10, 1990  Continue reading the main story Share This Page  Continue reading the main story  About the Archive  This is a digitized version of an article from The Times’s print archive, before the start of online publication in 1996. To preserve these articles as they originally appeared, The Times does not alter, edit or update them.  Occasionally the digitization process introduces transcription errors or other problems. Please send reports of such problems to archive_feedback@nytimes.com.  May 10, 1990, Page 00006 The New York Times Archives  HERE is a sampling of some of the better antiques and flea markets around the United States.  Two or Three Times a Year  BRIMFIELD Route 20, Brimfield, Mass. 01010; 413-245-3436. Second weekend of May and July, and the second weekend after Labor Day.  RENNINGER’S OUTDOOR EXTRAVAGANZA Noble Street, Kutztown, Pa.; 717-385-0104. Thursday, Friday and Saturday of the last weekend of April, June, September.  FARMINGTON ANTIQUES WEEKEND Farmington Polo Grounds, Town Farm Road, Farmington, Conn. 06032; 508-839-9735. Starting Wednesday before shows open; 203-677-7862. June 9-10 and Sept. 1-2.  Advertisement  Continue reading the main story  Monthly  ANN ARBOR ANTIQUES MARKET, P.O. Box 1512, Ann Arbor, Mich. 48106; 313-662-9453. May through October, third Sunday.  Continue reading the main story  Advertisement  Continue reading the main story  KANE COUNTY FLEA MARKET, Kane County Fairgrounds, P.O. Box 549, St. Charles, Ill. 60174; 708-377-2252. Year-round, first weekend.  THE METROLINA EXPO, 7100 Statesville Road, Charlotte, N.C. 28213; 704-596-4643. Year-round, first weekend of every month.  SPRINGFIELD ANTIQUE SHOW AND FLEA MARKET, Clark County Fairgrounds, Route 41, Springfield, Ohio, 45501; 513-325-0053. Year-round, third weekend.  Weekly  BAKERSFIELD SWAP-O-RAMA, 4501 Wible Road, Bakersfield, Calif. 93313; 805-831-9342. Saturday and Sunday.  LAMBERTVILLE ANTIQUE MARKET, Route 29, Lambertville, N.J. 08530. Weekend number: 609-397-0456. Weekday: 215-752-4485, between 5 and 7 P.M. Market on Saturday and Sunday.  ATLANTA FLEA MARKET AND ANTIQUE CENTER, 5360 Peachtree Industrial Boulevard, Chamblee, Ga. 30341; 404-458-0456. Friday, Saturday and Sunday.  A version of this list appears in print on May 10, 1990, on Page C00006 of the National edition with the headline: A Guide To Markets. Order Reprints| Today’s Paper|Subscribe  Continue reading the main story  What’s Next  Loading...  Go to Home Page »  Site Index The New York Times  Site Index Navigation  News  World  U.S.  Politics  N.Y.  Business  Tech  Science  Health  Sports  Education  Obituaries  Today’s Paper  Corrections  Opinion  Today’s Opinion  Op-Ed Columnists  Editorials  Op-Ed Contributors  Letters  Sunday Review  Video: Opinion  Arts  Today’s Arts  Art & Design  Books  Dance  Movies  Music  N.Y.C. Events Guide  Television  Theater  Video: Arts  Living  Automobiles  Crossword  Food  Education  Fashion & Style  Health  Jobs  Magazine  N.Y.C. Events Guide  Real Estate  T Magazine  Travel  Weddings & Celebrations  Listings & More  Reader Center  Classifieds  Tools & Services  N.Y.C. Events Guide  Multimedia  Photography  Video  NYT Store  Times Journeys  Subscribe  Manage My Account  NYTCo  Subscribe  Subscribe  Home Delivery  Digital Subscriptions  Crossword  Email Newsletters  Gift Subscriptions  Group Subscriptions  Education Rate  Mobile Applications  Replica Edition  Site Information Navigation  © 2019 The New York Times Company  Home  Search  Accessibility concerns? Email us at accessibility@nytimes.com. We would love to hear from you.  Contact Us  Work With Us  Advertise  Your Ad Choices  Privacy  Terms of Service  Terms of Sale  Site Information Navigation  Site Map  Help  Site Feedback  Subscriptions
@@
     [Trafilatura]    Possible Duplicate:  When should I use an em-dash, an en-dash, and a hyphen?  When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?  Possible Duplicate:  When should I use an em-dash, an en-dash, and a hyphen?  When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?  This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.  The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:  An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.  So when do you use an en-dash? Again from Mental Floss:  To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)  The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”  What about an em-dash? From here:  Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.  If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.
@@
     [Resiliparse]    1    Possible Duplicate:  When should I use an em-dash, an en-dash, and a hyphen?    When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?    marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32    This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.    0    The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:    An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.    So when do you use an en-dash? Again from Mental Floss:      1. To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)      2. The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”    What about an em-dash? From here:    Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.    If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.    Not the answer you’re looking for? Browse other questions tagged or ask your own question.
@@
     [WET file]    syntax - What’s the difference between - and -- in a phrase? - English Language & Usage Stack Exchange  Stack Exchange Network  Stack Exchange network consists of 175 Q&A communities including Stack Overflow, the largest, most trusted online community for developers to learn, share their knowledge, and build their careers.  Visit Stack Exchange  Log In Sign Up  current community  English Language & Usage  help chat  English Language & Usage Meta  your communities  Sign up or log in to customize your list.  more stack exchange communities  company blog  Tour Start here for a quick overview of the site  Help Center Detailed answers to any questions you might have  Meta Discuss the workings and policies of this site  About Us Learn more about Stack Overflow the company  Business Learn more about hiring developers or posting ads with us  By using our site, you acknowledge that you have read and understand our Cookie Policy, Privacy Policy, and our Terms of Service.  English Language & Usage Stack Exchange is a question and answer site for linguists, etymologists, and serious English language enthusiasts. Join them; it only takes a minute:  Sign up  Here’s how it works:  Anybody can ask a question  Anybody can answer  The best answers are voted up and rise to the top  Home  Questions  Tags  Users  Unanswered  What’s the difference between - and — in a phrase? [duplicate]  Ask Question  1  Possible Duplicate:  When should I use an em-dash, an en-dash, and a hyphen?  When do I put a - in a sentence? Is it a more powerful comma? With a bigger pause?  syntax dashes symbols  share|improve this question  edited Apr 13 ’17 at 12:38  Community  1  asked Jul 13 ’11 at 0:10  curiouscurious  123115  marked as duplicate by waiwai933, MrHen, user2683, Robusto, Thursagen Jul 13 ’11 at 0:32  This question has been asked before and already has an answer. If those answers do not fully address your question, please ask a new question.  add a comment |  1 Answer 1  active oldest votes  0  The dashes you described are known respectively as the en-dash and the em-dash. To describe the difference between their origins, Mental Floss writes:  An en dash (–) is bigger than a hyphen but shorter than an em dash (—). Th e names come from an obscure typographical measurement system, but the dashes have now taken on a life of their own in grammar. The em dash is the spork of English grammar: It ain’t particularly pretty, but you can use it for most anything. Em dashes can replace colons or sets of parentheses, or represent a sudden change in thought or tone.  So when do you use an en-dash? Again from Mental Floss:  To show numerical ranges, signifying “up to and including”—of dates, ages, pages, etc. (Example: “I read pages 7–22 last night.”)  The storied “compound adjective hyphen,” an event so rare in the English language that proofreaders shiver with excitement whenever they come across it. Basically “pro-American” gets a regular hyphen because “American” is only one word, whereas “pro–Falkland Islands” gets an en dash because “Falkland Islands” is two words. So, too phrases like “Civil War–era.”  What about an em-dash? From here:  Similar to an extended hyphen (-), an em dash is used to show a break in thought or a shift of tone.  If you’d like to read more about the differences between a hyphen (-), en-dash (–), and em-dash (—), see the blog post here which summarizes the above.  share|improve this answer  edited Aug 4 ’15 at 17:00  zwol  2,51911424  answered Jul 13 ’11 at 0:16  simchonasimchona  30.9k5112139  add a comment |  Not the answer you’re looking for? Browse other questions tagged syntax dashes symbols or ask your own question.  asked  7 years, 9 months ago  viewed  2,336 times  active  3 years, 8 months ago  Featured on Meta  Announcing the arrival of Valued Associate #679: Cesar Manara  Planned maintenance scheduled April 17/18, 2019 at 00:00UTC (8:00pm US/Eastern)  Linked  279  When should I use an em-dash, an en-dash, and a hyphen?  Related  279  When should I use an em-dash, an en-dash, and a hyphen?  4  When to use -, – and —?  6  What is the difference between `-` and `--`  5  Does “cost-benefit ratio” use a hyphen or an en-dash?  0  What kind of dash character should I use at the end of a famous saying to mark of the author?  -1  dash non-restrictive element in the middle of a sentence  1  em dash followed by a comma  0  what’s the difference between a hyphen, a dash and a minus sign?  0  Using comma to delimit the name of a group and its constituents?  1  I tend to overuse the hyphen as a pause, and would appreciate some feedback on this  Hot Network Questions  How could we fake a moon landing now?  Using audio cues to encourage good posture  Chinese Seal on silk painting - what does it mean?  How does the math work when buying airline miles?  What causes the direction of lightning flashes?  Maximum summed subsequences with non-adjacent items  Most bit efficient text communication method?  Significance of Cersei’s obsession with elephants?  Is it a good idea to use CNN to classify 1D signal?  What font is "z" in "z-score"?  How can I use the Python library networkx from Mathematica?  Fundamental Solution of the Pell Equation  Dating a Former Employee  Do I really need recursive chmod to restrict access to a folder?  Generate an RGB colour grid  How to Make a Beautiful Stacked 3D Plot  How to find all the available tools in mac terminal?  Has negative voting ever been officially implemented in elections, or seriously proposed, or even studied?  What is the meaning of the simile “quick as silk”?  Circuit to "zoom in" on mV fluctuations of a DC signal?  What are the out-of-universe reasons for the references to Toby Maguire-era Spider-Man in Into the Spider-Verse?  What does this Jacques Hadamard quote mean?  Can a new player join a group only when a new campaign starts?  Why wasn’t DOSKEY integrated with COMMAND.COM?  more hot questions  English Language & Usage  Tour  Help  Chat  Contact  Feedback  Mobile  Company  Stack Overflow  Stack Overflow Business  Developer Jobs  About  Press  Legal  Privacy Policy  Stack Exchange  Network  Technology  Life / Arts  Culture / Recreation  Science  Other  Stack Overflow  Server Fault  Super User  Web Applications  Ask Ubuntu  Webmasters  Game Development  TeX - LaTeX  Software Engineering  Unix & Linux  Ask Different (Apple)  WordPress Development  Geographic Information Systems  Electrical Engineering  Android Enthusiasts  Information Security  Database Administrators  Drupal Answers  SharePoint  User Experience  Mathematica  Salesforce  ExpressionEngine® Answers  Stack Overflow em Português  Blender  Network Engineering  Cryptography  Code Review  Magento  Software Recommendations  Signal Processing  Emacs  Raspberry Pi  Stack Overflow Ha pycckom  Programming Puzzles & Code Golf  Stack Overflow en español  Ethereum  Data Science  Arduino  Bitcoin  more (31)  Photography  Science Fiction & Fantasy  Graphic Design  Movies & TV  Music: Practice & Theory  Worldbuilding  Seasoned Advice (cooking)  Home Improvement  Personal Finance & Money  Academia  Law  more (15)  English Language & Usage  Skeptics  Mi Yodeya (Judaism)  Travel  Christianity  English Language Learners  Japanese Language  Arqade (gaming)  Bicycles  Role-playing Games  Anime & Manga  Puzzling  Motor Vehicle Maintenance & Repair  more (33)  MathOverflow  Mathematics  Cross Validated (stats)  Theoretical Computer Science  Physics  Chemistry  Biology  Computer Science  Philosophy  more (10)  Meta Stack Exchange  Stack Apps  API  Data  Blog  Facebook  Twitter  LinkedIn  site design / logo © 2019 Stack Exchange Inc; user contributions licensed under cc by-sa 3.0 with attribution required. rev 2019.4.18.33353  English Language & Usage Stack Exchange works best with JavaScript enabled
@@
MinHash is a locality-sensitive hashing technique used to group sets into collections based on their Jaccard similarity [28]. In the context of deduplicating text datasets, MinHash was first employed in Lee et al. [88] and then used in numerous other projects [121, 43]. We point readers to the main text of Lee et al. [88] and Appendix G.3.1 of Penedo et al. [121] for more details. The primary hyperparameters of note are the ngram-size, and the number of permutations used. Following Lee et al. [88], Penedo et al. [121], we use an ngram-size of 5 tokens and target a Jaccard similarity of 0.8. Departing from prior work, however, we modify the number of MinHash permutations used. Both Lee et al. [88] and Penedo et al. [121] use a total of 9,000 permutations, split into 450 buckets of 20 hashes each. We found this to be overly expensive and notice that similar Jaccard similarity plots can be attained with a much smaller number of permutations. For all of our ablations, we instead use a total of 1,395 permutations, split into 93 buckets of size 15. These hyperparameters were chose programmatically to mimic the Jaccard similarity plots as closely as possible, in an $\ell_{2}$ sense, with a fixed hash budget. See Figure 6 for more details.
@@
    In choosing a false positive rate for the ngram-based Bloom filter, it’s important to recognize that removal of a paragraph or document is dictated by having greater than a threshold fraction of the ngrams contained in the set. As an example, suppose we are given a paragraph of $N$ ngrams, where $S$ of them are already contained in the Bloom filter and we set threshold to $T$. Because Bloom filters do not allow false negatives, every one of the $S$ ngrams are marked (correctly) as contained, and $N-S$ of them could potentially be marked as a false positive. Indeed, of the $N-S$ of these ngrams, at least $TN-S$ of them would need to be marked as a false positive, each of which occurs independently with probability $\epsilon$. This is equivalent to $N-S$ Bernoulli random variables with parameter $\epsilon$, and can be bounded by a crude Hoeffding bound. In this particular case, the probability that a document or paragraph is falsely marked as a duplicate is bounded by: